I was working on my homework and this might be a stupid question but I'm only in 7th grade taking accelerated math. Anyway there was this problem 2s+12r and I'm suppose to restate each expression using distributive propery. How do I do that???? I answered 14xsr is the answer is that right???
Found 2 solutions by hulamc723, lkwtcnj:
Answer by hulamc723(15)

2s+12r
Using the distributive property means that you just need to find a common denominator and distribute it out
Meaning...
distribute it out
2(s+6r)
Now check the problems
(2*s)+(2*6r)=2s+12r
-mc

Answer by lkwtcnj(8)

It's not a stupid question! The distributive property lets you factor out common multiples in an expression. In this case, your only common factor would be 2 (since 2 and 12 are both divisible by 2). Just divide two out of the expression and place it outside:
2s + 12r = (2*1)s + (2 * 6)r = 2(s + 6r)
Good luck in your class!